### v1

# {tocentry} Vibrational Energy Dissipation in Non-Contact Single-Molecule Junctions Governed by Local Geometry and Electronic Structure

Abstract
The vibrational dynamics of adsorbate molecules in single-molecule junctions depend critically on the geometric structure and electronic interactions between molecule and substrate. Vibrations, excited mechanochemically or by external stimuli, dissipate energy into substrate electrons and phonons. Energy dissipation leads to the broadening of spectral lines, vibrational lifetimes, and the coupling between molecular and substrate phonons. It affects molecular manipulation, giving rise to nanoscale friction, and contributes to scanning probe and surface spectroscopy signals. We present an approach to disentangle adsorbate vibrational dynamics in non-contact junctions by employing density functional theory, machine learning, and non-adiabatic molecular dynamics. Focusing on the CO-functionalised Cu surfaces representing a single-molecule junction, a widely studied system in scanning probe and energy dissipation experiments, we reveal strong vibrational mode specificity governed by the interplay of electron-phonon and phonon-phonon coupling. Electron-phonon relaxation rates vary by two orders of magnitude between modes and sensitively depend on the tip-substrate geometry. We find evidence of a weak non-additive effect between both energy dissipation channels, where electron-phonon coupling enhances phonon-phonon coupling. Our predicted vibrational lifetimes agree with infrared spectroscopy and helium scattering experiments. Finally, we outline how our findings can inform and enhance scanning probe experiments.
keywords:
mechanochemistry, single-molecule microscopy, electron-phonon coupling, anharmonic phonon coupling, density functional theory, vibrational energy dissipation1 Introduction
The vibrational dynamics of atoms and molecules adsorbed on a surface provide valuable insights into the interactions and energy exchange between adsorbate and substrate. These dynamics are governed by interactions between adsorbates and the surface and quantum-mechanical energy dissipation mechanisms,1, 2 Energy dissipation mechanisms in non-contact single-molecule junctions significantly affect measured signals in atomic force microscopy (AFM), lateral force microscopy (LFM),3, 4 scanning tunnelling microscopy (STM),5 inelastic electron tunnelling spectroscopy (IETS), and infrared spectroscopy experiments.6, 7 In such experiments, a key challenge is to disentangle geometric structure, dynamics, and energy dissipation mechanisms in the measured signal: For example, AFM-induced hopping of atoms and molecules leads to measurable energy dissipation,8, 9, 10, 3 yet the underlying dissipation pathways remain unclear. Helium atom scattering (HAS) experiments of CO adsorbed on Cu surfaces have found a strong dependence of vibration frequencies on both the surface facet and the presence of defects.11
It is well-established that the surface geometry plays a critical role in shaping molecule-surface interactions and energy transfer processes: Previous studies have shown the effect of defects on static interfacial charge transfer,12 while adsorption site and facet affect the vibrational properties of adsorbates.13 Establishing an understanding of how surface geometry alters different energy dissipation channels and affects the vibrational dynamics of molecules adsorbed on surfaces remains an important research area.14
Figure 1 shows a schematic setup of an idealised single-molecule junction as it would be present in an AFM experiment. If vibrations of the molecule in the junction are excited, for example, through the motion of the tip, energy dissipation away from the molecule can occur by excitation of phonons or electron-hole-pair (EHP) excitations in the metal tip or substrate. The former is due to anharmonicity that induces phonon-phonon coupling between adsorbate and substrate.2 The latter is due to coupling between electronic and nuclear motions, commonly referred to as electron-phonon coupling.6 Intramolecular vibrational relaxation and radiative processes, such as infrared emission, also contribute to energy dissipation, though their role is typically minor for small adsorbates.15, 16 However, for larger adsorbates, intramolecular vibrational relaxation can become a significant dissipation channel.17 Consequently, it is of great interest to explore how the geometric structure of a surface or junction affects the mechanical and electronic interactions with the adsorbate, which govern the relative contributions of electron-phonon and phonon-phonon coupling to the vibrational dynamics. Gaining control over these energy dissipation channels requires the ability to identify their distinct signatures and disentangle their interplay—an essential step towards rational design of surface functionality at the atomic scale.
In this work, we study competing energy dissipation mechanisms and their effect on the vibrational dynamics of adsorbate molecules for the exemplary case of CO-decorated tip-surface copper junctions. This system has been extensively characterised by experiment18, 6, 19 and is widely used in atomic-scale resolution scanning probe measurements.20 Various previous theoretical studies of energy dissipation for CO on different, idealised Cu surfaces21, 22, 23 have been reported, but no clear analysis of the role of the local adsorption geometry exists to date. A specific challenge in this system for electronic structure theory is the correct description of the potential energy surface of CO on Cu(111).24, 25, 26 We address this by using the state-of-the-art screened hybrid functional HSE0627 in combination with the MBD-NL van der Waals (vdW) correction28 to correctly describe the adsorption site and the potential energy surface. From a first-principles perspective, we describe the details and main energy dissipation mechanisms, namely electron-phonon and phonon-phonon coupling, which govern the dynamics at the atomic scale.2 We find that the surface geometry plays a significant role in these mechanisms, with flatter surfaces exhibiting stronger electron-phonon and phonon-phonon coupling. Additionally, we observe a strong mode dependence of electron-phonon coupling, with relaxation rates differing by two orders of magnitude. The molecular dynamics with electronic friction approach29, 21 allows us to study the interplay of the two dissipation channels and their concerted effects on the relaxation dynamics. We find that the two effects display weak non-additivity, suggesting that subtle dynamical steering may arise from one or both mechanisms. Our results are in good agreement with existing experiments and support the interpretation of single-molecule dynamics in non-contact junctions.
2 Results
2.1 Models of tip geometry in non-contact junctions
As shown in Figure 1, we can consider two interaction regimes for a non-contact CO tip-molecule-surface junction: In one case, the molecule is attached to the surface, in the other case, it is attached to the tip. The molecule can be excited by a variety of processes: The motion of the tip will displace the molecule with respect to the corrugated potential energy landscape of the underlying substrate.30 The mechanochemical forces the molecule experiences in the process will excite its vibrations. The molecular vibrations can absorb light,6 supported by field enhancement in the junction.31 A potential bias and electron tunnelling can lead to current-induced forces.3 In all cases, molecular vibrations will be excited, leading to energy dissipation into EHPs and phonons in the metal. The case where the molecule is attached to the surface is not too different from the scenario in the absence of the tip. Due to multiple Angstroms of distance between the tip and the molecule, mechanochemical forces on the molecule will be weak. Other perturbation scenarios have previously been studied.23, 22 For simplicity, we consider here only surfaces in both the slab and slab–adatom systems that correspond to the Cu(111) facet.
We turn our attention to the second case, where the molecule is attached to the tip. Regardless of the origin of the vibrational excitation, energy dissipation into the underlying substrate in this scenario will be almost negligible, as evidenced by calculations for CO on Cu in Section S1 of the Supporting Information. Therefore, molecular vibrations will dominantly dissipate energy into EHPs and phonons in the tip. This motivates our model design: We consider four different possible “tip” geometries: CO on a sharp tip (sharp tip), CO on a blunt tip (blunt tip), CO on an adatom on a surface slab (slab adatom), and CO on a slab (slab), which are depicted in Figure 1c and Figure S3 in the Supporting Information. The underlying surface, in the case of lateral motion of the tip, only serves as a corrugated energy landscape that triggers molecular vibrations4 and therefore does not have to be explicitly considered in the atomistic model. The “tip” models are therefore represented by two-dimensional periodic surface slab geometries. Further details on the models and the computational methodology are provided in the Methods section and Section S1 of the Supporting Information.
2.2 Vibrational mode analysis
CO adsorbed on a flat metal surface has four distinct non-degenerate vibrational modes: frustrated translation (FT), frustrated rotation (FR), surface-adsorbate stretch (SA), and the internal stretch of the CO bond (IS), as illustrated in Figure 2. The FT and FR modes consist of two degenerate modes each.
We determine vibrational energies from second-order force constants (i.e. the Hessian) for all atoms in the molecule and the top layers of the four surface models. Simulations include ten layers of freely movable Cu atoms, spanning hundreds of degrees of freedom. Energies, forces, and second-order force constants are evaluated with a MACE graph neural network model32 trained on first-principles data (see Methods). The MACE potential accurately reproduces the underlying HSE06 data, correctly predicting the adsorption sites: the top site for CO on a (111) surface slab of copper and the hollow site for the adatom in the CO-adatom arrangement. (see Section S2 of the Supporting Information). The calculated vibrational frequencies for CO on the Cu(111) slab show good agreement with experimental data, as presented in Table 1. Interestingly, the stiffness of low-energy modes increases as the tip becomes blunter, as indicated in Table 1. Moreover, there is a jump in stiffness when transitioning from CO adsorbed on an adatom to CO adsorbed directly on the surface. HAS experiments estimated the energy of the FT modes on a flat Cu(111) surface to be , while the energy dropped to at step edges of the Cu(211) and Cu(511) surfaces.11 This trend is in good agreement with our finding that more tapered (or corrugated) surface geometries exhibit lower vibrational energies for the FT mode.
vibrational energies / meV | ||||
FT | FR | SA | IS | |
sharp tip | 2.74 | 31.38 | 49.06 | 271.48 |
blunt tip | 2.89 | 32.28 | 50.35 | 270.86 |
slab adatom | 3.17 | 31.74 | 49.13 | 272.26 |
slab top | 5.29 | 35.92 | 47.97 | 264.51 |
slab top exp.11, 7 | 4.07 | 36.47 | 41.41 | 257.63 |
The vibrational density of states (VDOS) (see Figure 3) provides an overview of all vibrational modes. We differentiate between modes localised dominantly at the molecule (orange in Figure 3), where the majority of the kinetic energy is contained in the CO and substrate modes (black in Figure 3), where most of the energy is in the motion of the Cu atoms. The FR, SA, and IS vibrational modes of CO exhibit higher frequencies compared to most copper substrate modes, which typically range between approximately and . We note that the VDOS shown here is normalised by the number of atoms. The substrate vibrational states are more localised in energy (or frequency) for slab and slab-adatom surfaces and more delocalised for the sharp and blunt tips. For sharp and blunt tips, a significant portion of the surrounding space is vacuum, whereas this space is occupied by atoms in the slab and slab-adatom surfaces. The presence of more atoms in these flatter configurations results in a greater number of substrate vibrational states per volume. This facilitates stronger overall coupling between the CO and substrate vibrational modes, provided that the CO and Cu degrees of freedom are sufficiently coupled. As a rule of thumb, if the off-diagonal elements of the Hessian (for CO-Cu coupling) are of a similar magnitude, we can expect strong coupling. As shown in Figure S4 in the Supporting Information, this is indeed the case. In the case of the sharp and blunt tip, we find a strong coupling for the CO to a small number of Cu degrees of freedom. The slab adatom and slab geometries exhibit weaker coupling but a larger number of states. When discussing phonon-phonon coupling, we will return to this observation.
2.3 Phonon-phonon coupling
The atomic motion of the CO molecule is coupled to the atomic motion of the Cu tip through phonon-phonon interactions and to EHP through electron-phonon coupling.2 Even at low temperatures, phonon-phonon coupling plays a significant role in the dynamics of molecular vibrations. In our study, we use a temperature of , a condition commonly employed in experiments such as AFM measurements with CO-decorated tips. Phonon-phonon coupling is determined by analysing the dynamics of a CO molecule on surface geometries where approximately 500 Cu atoms are free to move per unit cell, allowing energy transfer between CO modes and substrate modes. To determine coupling strengths, we use vibrational analysis of molecular dynamics trajectories in thermodynamic equilibrium.33 We first thermalise each system at a temperature of and then allow it to evolve in thermodynamic equilibrium over time. Over time, the velocity time series of the atoms becomes decorrelated, as the atoms transfer mechanical energy between each other. The rate of this decorrelation can be found in the cross-correlation function. For this analysis, we employ normal mode decomposition to project the velocity time series onto the basis defined by the vibrational eigenvectors. For the velocity time series of two modes and the velocity cross-correlation function is as follows:
The Fourier transform of the velocity cross-correlation function yields the cross-spectrum:
The cross-spectrum provides valuable insights into vibrational lifetimes and relative coupling strengths. Let us first discuss the coupling strength, which can be determined by analysing the off-diagonal elements . These exhibit two Lorentzian-shaped peaks at the characteristic frequencies of modes and , with their heights reflecting the coupling strength between these modes (see Section S5 of the Supporting Information). To further understand this interaction, we examine the coupling density , defined as the sum of the cross spectra for all modes correlated with the mode of interest :
A broad peak in the coupling density at a given frequency means that modes at this frequency are correlated with the mode of interest, , indicating strong vibrational coupling. Our analysis reveals that the coupling is strongest with adjacent substrate vibrational modes, as shown in Figure 4 and Figures S18, S19, S20, and S21 in the Supporting Information. This effect is less pronounced in case of energetically decoupled SA and IS modes. The coupling between different CO modes is small in comparison, indicating that intramolecular vibrational redistribution plays only a minor role. Notably, the FR and IS modes exhibit the strongest coupling.
The geometry of the surface has a substantial impact on the coupling density. Among the different geometries, the sharp tip exhibits the weakest coupling, while the slab geometry demonstrates the strongest coupling density. This observation can be attributed to the VDOS discussed earlier. As surface geometry becomes more atomically flat, the VDOS increases, offering more vibrational states to accommodate the energy of CO modes and thereby enhancing phonon-phonon coupling.
The diagonal terms of the cross-correlation function contain the power spectrum. The power spectrum exhibits a single Lorentzian-shaped peak at the characteristic frequency of the mode , where the full width at half maximum is inversely proportional to the lifetime of mode .
As shown in Table 2, the overall lifetimes tend to decrease as the surface becomes more atomically flat. Interestingly, this trend does not apply uniformly to all modes. The FT mode shows the strongest dependence on surface geometry, while this dependence is less pronounced for the FR and SA modes. The lifetime of the IS mode remains largely unaffected by changes in surface shape. Overall, the influence of the surface geometry is strongest in low-energy modes and becomes weaker as the energy of the mode increases. The largest phononic lifetime (FT on sharp tip model) is one order of magnitude larger than the smallest lifetime (FT on slab model).
Lifetime / ps | ||||
FT | FR | SA | IS | |
sharp tip | 142.1 | 69.0 | 56.2 | 48.5 |
blunt tip | 110.9 | 67.0 | 31.9 | 65.4 |
slab adatom | 59.8 | 65.3 | 53.7 | 46.3 |
slab | 19.3 | 24.6 | 26.8 | 53.7 |
2.4 Electron-phonon coupling
Electron-phonon coupling is a significant factor in vibrational energy dissipation on metal surfaces.34, 2 Due to the absence of a band gap, the Born-Oppenheimer approximation does not strictly hold, and the motion of adsorbate atoms can induce EHP excitations that lead to nonadiabatic effects, sometimes also called electronic friction effects. In the weak perturbation regime, e.g. where adsorbate motion is triggered mechanochemically or through vibrational excitation, this effect leads to nonadiabatic energy dissipation. To describe the impact of these excitations on dynamics, different simulation methods have been developed. Time-dependent perturbation theory has emerged as a feasible approach for extended systems. This method enables the calculation of the electron-phonon-induced vibrational lifetimes (see Methods),35 which we discuss in the following.
We evaluate electron-phonon coupling at a temperature of , typical for non-contact single-molecule junction experiments. At such low temperatures, electron-phonon coupling dominates over phonon-phonon coupling in metals.36, 37 At this temperature, first-order electron-phonon coupling is a dominant contribution to vibrational lifetimes and higher-order electron-phonon coupling effects can be neglected . 38 Its strength is sensitive to the electronic states that interact with nuclear motion, which we characterize by the electronic density of states (DOS) projected onto the copper atom bonded to CO (PDOS). The PDOS is highly sensitive to the geometry of the system. Since the energy to excite electrons is supplied by vibrational states with frequency , only a small portion of the DOS near the Fermi level ()—relative to the energy scale of electronic states—is accessible. Therefore, analysing the DOS at the Fermi level is instructive for understanding the strength of electron-phonon coupling. As shown in Figure 5, the PDOS of the sharp and blunt tips exhibits states localised in energy. The magnitude of the PDOS close to the Fermi level is small for the sharp tip, while the blunt tip exhibits energetically localised spikes. In contrast, the slab and slab-adatom systems have a PDOS that is delocalized in energy. The slab PDOS exhibits a larger magnitude compared to that of the adatom on the slab. Since a larger magnitude of the PDOS at and close to the Fermi level allows for stronger electron-phonon coupling, we expect that electron-phonon coupling will be stronger at flatter surfaces.
The strength of electron-phonon coupling is reflected in the vibrational lifetimes, where shorter lifetimes indicate stronger coupling (see Table 4). The FT and SA modes experience relatively low electron-phonon coupling (resulting in longer lifetimes), whereas FR and IS modes exhibit electron-phonon coupling approximately an order of magnitude higher (with lifetimes an order of magnitude shorter) compared to FT and SA. Previous studies of CO on different Cu facets23, 22, 21 have reported similar trends of mode dependence, as we show in Table 3. Our results for the slab are in good agreement with literature values. In the case of the slab, there is an ongoing debate about whether the FR or the IS decays faster. Indeed, the lifetimes reported here and in other literature are very similar for these modes. However, as the tip becomes more corrugated, the IS mode clearly becomes the fastest dissipation mode, and the lifetime of the FR mode increases. As shown in Table 3, vibrational lifetimes are significantly longer on sharper tips, with CO on a flat surface exhibiting the shortest lifetimes. This trend aligns with the reduced magnitude of the PDOS close to the Fermi level on more corrugated surfaces.
Lifetime / ps | ||||
FT | FR | SA | IS | |
sharp tip | 185.9 | 21.2 | 80.0 | 2.8 |
blunt tip | 99.1 | 6.0 | 39.0 | 3.3 |
slab adatom | 110.8 | 10.8 | 52.9 | 4.9 |
slab (Cu(111)) | 47.9 | 3.6 | 12.5 | 1.9 |
CO on Cu(111)23 | 84.4 | 2.1 | 7.8 | 2.2 |
CO on Cu(110)22 | 76.9 | 2.7 | 10.8 | 2.2 |
CO on Cu(100)21 | 125.0 | 2.6 | 20.0 | 3.2 |
2.5 Electron-phonon and phonon-phonon coupling
The experimentally measured lifetime is governed by both electron-phonon and phonon-phonon coupling. To capture this combined effect, we employ Langevin dynamics, specifically molecular dynamics with electronic friction (MDEF).29 We use a classical approximation, neglecting the zero-point energy of the phonons. This approach has been successfully applied to the dynamics of atoms39, 40 and molecules36, 37 at metal surfaces. In MDEF, electron-phonon coupling is treated as a friction force represented by the electronic friction tensor (see Equation 9). The electronic friction tensor is calculated in the Markovian quasi-static limit. Here, the frictional force experienced by the nuclei depends only on their instantaneous velocities, without memory (non-Markovian) effects from past states. This assumes that the electronic structure adjusts instantaneously to the nuclear configuration. In the quasi-static limit, the electronic friction tensor is computed for a given nuclear geometry without explicitly propagating electronic dynamics. This approach is common for molecular dynamics on metals.35, 41 We extract the total vibrational lifetime from phonon-phonon and electron-phonon contributions from MDEF simulations, by utilising two complementary methods. Equilibrium correlation analysis of the velocity cross-correlation function of MDEF trajectories in thermodynamic equilibrium allows determining lifetimes, as we have done for the phonon-phonon coupling. Kinetic energy decay predicts lifetimes by evaluating the explicit non-equilibrium decay of the kinetic energy of vibrational modes after selectively energizing a given mode. For both approaches, we employ normal mode decomposition to project the velocity time series onto the basis defined by the vibrational eigenvectors. Additionally, lifetime can be estimated from the independent phonon-phonon and electron-phonon lifetimes presented in Tables 3 and 2. Additive relaxation rates estimate the total lifetime by summing the phonon-phonon and electron-phonon relaxation rates predicted independently in previous sections.
Equilibrium correlation analysis. Table 4 shows the vibrational lifetimes from equilibrium correlation analysis of the MDEF simulation results. While the IS and FR lifetimes are dominated by the electron-phonon contribution, the FT mode is dominated by the phonon-phonon contribution. The two effects are somewhat equal in contribution to the SA mode lifetime. Lifetimes for the FR and IS modes are in good agreement with infrared spectroscopy experiment. However, the experimental lifetime of the SA mode is , while the calculated lifetime is for the Cu(111) surface. A similar discrepancy has been reported in literature and was attributed to inhomogeneous broadening:6 Experiments conducted on rougher surfaces show a greater sensitivity of the mode amplitude and linewidth of the SA mode (and the amplitude of the FR mode) than for the IS mode. We note in passing that the literature lifetimes from infrared spectroscopy experiments should be considered as lower bounds (marked by asterisks (*) in Table 4). This is because the observed linewidths are artificially broadened due to limitations of the measurement setup. Since lifetime is inversely related to linewidth, this broadening leads to a systematic underestimation of the lifetimes. Additionally, helium scattering experiments have measured the lifetime of the FT mode of CO on Cu(100) to ,42 which is in the same order of magnitude as the lifetime for the FT mode of CO on a Cu(111) slab reported here.
Lifetime / ps | ||||
FT | FR | SA | IS | |
sharp tip | 40.4 | 24.1 | 21.1 | 5.2 |
blunt tip | 30.0 | 4.1 | 15.2 | 9.0 |
slab adatom | 21.5 | 8.5 | 28.4 | 6.3 |
slab | 19.8 | 3.6 | 20.7 | 3.6 |
CO on Cu(111)6 | - | 2.5* | 2.5* | 2.5* |
CO on Cu(100)6 | - | 2.5* | 1.9* | 0.7* |
CO on Cu(100)43 | - | 8 | - | - |
Kinetic energy decay. We can compare lifetimes determined through equilibrium correlation analysis with those explicitly calculated from non-equilibrium simulations (see Method section). In these simulations, we excite vibrations with a specific amount of energy and monitor the decay of the kinetic energy (Figure 6). To determine the lifetimes, we assume that the vibrational modes behave like damped harmonic oscillators, where the envelope of the kinetic energy is assumed to exhibit an exponential decay. Overall, we observe that the kinetic energies of the CO vibrational modes indeed show an exponential decay. A comparison between Table 4 and Figure 6 demonstrates that the non-equilibrium lifetimes agree well with those from equilibrium correlation analysis.
Additive relaxation rates. The total relaxation rate , or the inverse of the lifetime , can be approximated by summing the contributions from electron-phonon (Table 3) and phonon-phonon coupling (Table 2).
Figure 7a compares the three methods of determining the total lifetime (or relaxation rate). Within the estimated uncertainty of our simulations, the results are in good agreement. All three methods yield lifetimes that exhibit similar trends across the different systems and modes. Flatter surface geometries lead to overall larger relaxation rates and smaller lifetimes. Relaxation rates exhibit a strong mode dependence, with the largest rate being an order of magnitude greater than the smallest. The numerical values in Figure 7a can be found in Section S3 of the Supporting Information.
Contributions to the relaxation rates. Let us now disentangle the phonon-phonon and electron-phonon contributions to the relaxation rates in dynamic simulations. Figure 7b shows the two components that we decompose using the kinetic energy decay method. The phonon-phonon relaxation rates are determined using a similar procedure to that for the total relaxation rate, but exclusively with MD and without electronic friction. The phonon-phonon relaxation rates are obtained by monitoring the decay in kinetic energy during MD simulations (similar to total rates but without electronic friction). The electron-phonon relaxation rate is extracted from the MDEF trajectories by integrating the electronic friction and random force components (see equation 9). We note that the random force component integrates to zero in equilibrium.
Figure 7b shows that overall, the relaxation rate (and consequently the lifetimes) of the FT and SA modes are primarily dominated by phonon-phonon coupling, whereas the FR and IS modes are predominantly influenced by electron-phonon coupling. Interestingly, the sum of phonon-phonon and electron-phonon contributions does not exactly match the total relaxation rate. Overall, we observe a trend to underestimate the total rate if individual components are summed. While this trend to underestimate the total relaxation rate is within the uncertainty of our simulation, it does hint towards some level of non-additivity and a deviation from equation 5 in the MDEF dynamics. Moreover, we have assumed a constant electronic friction tensor. In reality, the friction tensor is configuration dependent and changes as the geometry is displaced along different vibration modes. This effect will likely further add to non-additivity. Electron-phonon and phonon-phonon relaxation rates may not be fully independent, as shown in Figures S18, S19, S20, and S21 in the Supporting Information. There, we find that the phonon-phonon coupling between the CO and the substrate vibrational modes increases if electron-phonon coupling is included in the equilibrium dynamics simulations.
Discussion
Vibrational modes and phonon-phonon coupling. Remarkably, all phononic lifetimes fall within the same order of magnitude. In contrast, electron-phonon lifetimes vary by two orders of magnitude (IS vs. FT), highlighting a much stronger mode dependence compared to phonon-phonon interactions. Additionally, despite the significantly higher vibrational energy of the IS mode, its lifetime remains comparable to other modes, suggesting a surprising degree of energy independence in phonon-phonon coupling. To understand this behaviour, we analyze the VDOS and coupling density. In the slab and slab-adatom systems, energetically more localized substrate vibrational states lead to reduced phonon-phonon coupling compared to the sharp and blunt tips. This is counteracted by a larger number of vibrational states per volume (approximately three times as many) in the slab and slab-adatom systems. The FT mode has an energy comparable to the lowest substrate vibrational modes, while the FR mode is close in energy to the largest substrate modes. Coupling of the molecular vibrations is most efficient to the substrate vibrations with similar energy, as shown in Figure 4 and Supporting Information Figures S18, S19, S20, and S21 (Section S5). We can understand this in the following way: Energy conservation mandates that any newly excited vibrational mode arises from the creation or annihilation of other modes, following . As the number of participating modes increases, the efficiency of the process decreases. Because the FT and FR modes overlap energetically with the substrate modes, they can access efficient processes. However, individual substrate vibrations are strongly dependent on the surface shape. As a result, the phononic lifetimes of the FT and FR modes depend strongly on the surface shape, as shown in Table 2. The SA and IS modes are energetically relatively decoupled from the substrate modes compared to FT and FR. Therefore, their coupling of the substrate modes involves high- processes. Supporting Information Figures S20 and S21 show that the SA and IS modes couple equally to all surface modes. Consequently, the influence of surface geometry on specific surface vibrations is effectively averaged out, as shown in Table 2.
Electron-phonon coupling. We find small electron-phonon lifetimes for the FR and IS modes and large lifetimes for the FT and SA modes, in good agreement with experimental and theoretical literature.6, 21, 23, 22 For each tip geometry, the electron-phonon lifetimes follow the trend of the vibrational frequencies of the respective modes, except for the FR mode. We can understand this behavior if we assume that the rate of EHP excitation is independent of the energies of the initial and final states. This corresponds to the wide-band, constant-coupling approximation, where one assumes a constant electronic DOS. In such a case, relaxation rates scale quadratically with the vibration energy. A similar explanation has been proposed by Persson.23 As the vibrational energy increases, the available energy window for electronic transitions broadens, enabling more excitation pathways and thereby decreasing the lifetime. The vibrational energy of the IS mode is two orders of magnitude higher than that of the FT mode, consistent with the two-order-of-magnitude difference in their lifetimes. The FR mode deviates from this trend, with lifetimes comparable to the IS mode for the slab geometry. This can be explained by the large electron-phonon coupling reported by STM vibrational spectroscopy experiments for this mode.44 Therefore, the EHP rate cannot be assumed to be independent of the involved energy states in this case.
Moving from the slab model to the sharp tip model, the lifetimes become larger. As Figure 5 shows, the magnitude of the electronic PDOS becomes smaller with more corrugated tips. This leaves fewer electronic states available for electronic excitations. Consequently, electron-phonon coupling becomes less efficient, and the lifetime increases. However, the IS mode deviates from this trend. The available vibrational energy of the IS mode is approximately , which is one order of magnitude larger than the energy of the SA and FR modes. The large vibrational energy of the IS mode increases the number of accessible electronic states for excitation. Even for the sharp tip, where the PDOS is highly localized, this energy range enables excitations between many more electronic states. Since the other modes carry significantly less energy, they are more strongly affected by the surface geometry: The lifetime of the FT mode increases by one order of magnitude between the slab and the sharp tip, making it two orders of magnitude longer than that of the IS mode.
Combined energy dissipation mechanisms. We have determined total relaxation rates using three methods: additive relaxation rates, equilibrium correlation analysis, and kinetic energy decay. Despite their significant conceptual difference, all three methods predict lifetimes that exhibit similar trends across the different systems and modes (see Figure 7a). This demonstrates the robustness of our analysis.
In Figure 7b, we have also observed that the sum of phonon-phonon and electron-phonon relaxation rates from dynamic simulations tends to be smaller than the MDEF relaxation rates. While this difference lies within the uncertainties and is therefore strictly not significant, it is in line with a comparison of the coupling densities with and without electron-phonon coupling: Coupling densities become larger if electron-phonon coupling is included in the simulations, indicating a weak non-additive effect where electron-phonon-mediated phonon-phonon coupling increases phononic relaxation. However, the approximations in our approach likely underestimate this non-additivity. We assume a constant electronic friction tensor, whereas in reality, it is configuration-dependent. As the geometry is displaced along different vibration modes, the strength of electron-phonon coupling changes, likely enhancing the degree of non-additivity.
The observation from Figure 7b that the relaxation rate of the FR (and IS) mode is dominated by electron-phonon coupling while the SA mode is not, is in good agreement with experiment. Asymmetric Fano line shapes have been reported for the FR mode, while symmetric Lorentzian line shapes have been measured for the SA mode.6 Fano line shapes typically indicate that a discrete vibrational mode interacts with a continuum of electronic states, indicating significant electron-phonon coupling. Conversely, a Lorentzian line shape suggests that phonon-phonon coupling dominates the relaxation of the vibration mode.45
However, it is important to note that the presence of electron-phonon coupling does not necessarily lead to a Fano line shape. According to Langreth,45 the line shape of a vibrational mode in the presence of EHP excitations is governed by the dynamic dipole , with the asymmetry depending on the imaginary part, :
Here and are the ionic and electronic dipole, respectively. The electronic dipole consists of a real and an imaginary part. The imaginary part emerges due to the creation of EHPs as adsorbate electrons and holes tunnel into the substrate. Since this tunneling process is not instantaneous, the oscillations of the electronic dipole get out of phase with the driving electric field.
The imaginary part of equation 6 becomes large—giving rise to an asymmetric line shape—if the imaginary electronic dipole is large compared to the ionic dipole . This occurs for modes with a dipole oriented parallel to the substrate—such as the FR mode—where screening by mirror charges significantly reduces the effective dipole strength. By contrast, modes with atomic motion perpendicular to the substrate, such as the SA and IS modes, exhibit a large ionic dipole. Subsequently, asymmetric line shapes would not be expected here, even if electron-phonon coupling is the dominant mechanism.
Perspective and potential experimental validation. Finally, equilibrium correlation analysis and kinetic energy decay yield lifetimes in good agreement, which demonstrates the validity of the former. Equilibrium correlation analysis can yield important information, such as coupling densities and phononic spectral functions. Our method for disentangling electron-phonon and phonon-phonon contributions to the vibrational dynamics of adsorbed molecules enables a better understanding of electronic and vibrational mass effects, as well as effects due to the local bonding environment. It can improve the interpretability of vibrational line widths and energy dissipation observed in AFM, IETS, Raman, or infrared spectroscopy. If the relaxation of a vibrational mode is primarily governed by electron-phonon coupling, adjusting electronic properties, such as the tip material, functionalization, or the bias voltage, can enhance signal detection and reveal details about the coupling. If phonon-phonon coupling dominates, tuning temperature or using a different isotope (14C or 13C instead of 12C in CO, for instance) can shift vibrational frequencies and affect the coupling. In scanning probe experiments and single-molecule manipulation, where vibrational excitation can induce molecular motion or reactions, understanding electron-phonon and phonon-phonon relaxation supports design conditions that favor specific molecular transformations. Similarly, in catalysis, targeting specific vibrational modes can selectively trigger reactions, enhancing the control and efficiency of chemical processes.
3 Conclusion
The vibrational dynamics of CO-decorated copper tips that range from atomically flat to maximally sharp exhibit pronounced vibrational mode specificity, driven by the interplay of electron-phonon and phonon-phonon coupling, as well as significant geometric effects. Lifetimes due to electron-phonon coupling differ by two orders of magnitude between modes due to vibrational energies, the symmetry of the mode, and the DOS. Overall, CO vibrations of flatter tip geometries exhibit stronger electron-phonon relaxation, due to the increased energetic delocalization of the local electronic DOS. However, the high-energy IS mode does not show this trend due to the large energy window to excite electrons. Phonon-phonon lifetimes also decrease as the surface becomes flatter, although the IS mode does not follow this trend. The FT, FR, and SA modes are coupled to the substrate vibrations, leading to significant geometry dependence, while the IS mode is decoupled from the substrate vibrations. The sum of individual electron–phonon and phonon–phonon relaxation rates obtained from dynamic simulations tends to underestimate the total relaxation rate computed by the same method. This weak non-additivity arises from electron-phonon-mediated phonon-phonon coupling, which enhances phonon-phonon interactions. Finally, the lifetimes obtained from equilibrium correlation analysis and kinetic energy decay are in good agreement and also agree well with experiment.
Future work will explore the role of non-diagonal electronic friction on the mode selectivity and surface dependence. Incorporating a fully configuration-dependent electronic friction tensor will enable a more accurate investigation of non-additivity.46 Finally, extending the current framework to account for zero-point energy contributions will further improve the fidelity of atomic-scale friction models.
4 Methods
4.1 Modelling the PES of CO on Cu(111)
Due to the many-electron self-interaction error, the energies of the states responsible for bonding the CO molecule to the Cu surface are misrepresented. This results in an incorrect preference for the fcc-hollow site over the top site, which experimental studies have shown to be the correct adsorption site.47 Generalized Gradient Approximation (GGA) functionals fail to accurately describe CO adsorption on Cu and other metal surfaces.24, 25, 26 The Random Phase Approximation (RPA) has been proposed as a method to yield qualitatively correct adsorption sites, while the PBE0 functional shows a slight preference for the correct adsorption site.48 However, RPA calculations are unfeasible due to their high computational cost, while PBE0 is not ideal for modelling bulk metals since it tends to overestimate the delocalization of electrons, which can lead to an inaccurate description of metallic systems, particularly their electronic structure and density of states near the Fermi level.49, 50 In our work, we use the HSE06 screened hybrid functional,27 which treats short-range interactions with exact exchange and reduces the amount of exact exchange at long range. This helps mitigate some of the issues, such as over-delocalization of electrons and poor convergence in calculations.51 We combine this function with the MBD-NL long-range dispersion correction28—a combination that was previously shown to provide accurate adsorption structure and energetics.52 With this combination, we predict an accurate PES (see Table S1 in the Supporting Information). Calculations are performed using the FHI-aims quantum chemistry code53 with periodic boundary conditions and a repeated slab approach, including a dipole correction. The sharp and blunt tip geometries are placed on a four-layer slab for better convergence. Convergence tests and further details are shown in Section S1 of the Supporting Information.
4.2 Interatomic potential
We use the MACE equivariant neural network potential32 and train it at the HSE06+MBD-NL level of theory, which accurately describes the potential energy surface of CO on Cu(111) as well as CO adsorbed on an adatom. Using a training set of 621 data points, the potential achieves an energy error of /atom and a force error of , which was determined using k-fold cross-validation, where we split our dataset into five equally sized segments. The training data contains structures of all four surface geometries (sharp tip, blunt tip, slab adatom, slab). Details regarding hyperparameters, training/test splits, and validation during dynamics simulations are shown in Section S2 in the Supporting Information.
4.3 Vibration calculatons
Second-order force constants are calculated using finite displacements of . Energies and forces are determined using the MACE potential. During the vibration calculation, the bottom four layers of our systems are frozen. The second-order force constants are calculated for the top 10 mobile Cu layers as well as the CO molecule (approximately 500 atoms).
4.4 Description of electron-phonon coupling
In first-order perturbation theory, we can express the elements of the electronic friction tensor associated with atomic degrees of freedom and as follows:35, 21
The electronic friction tensor associated with energy of a given vibration mode is calculated by summing over all possible excitations between effective single particle Kohn–Sham states and . The indices and indicate the initial and final state, respectively, and is the associated wave vector. The terms in the sum are given by the nonadiabatic coupling matrix elements (first row of Equation 7), the occupation factors and of the initial and final states (second row), and the conservation of energy (third row). The term accounts for the perturbing frequency of the vibration mode of interest. We calculate electronic friction with the FHI-aims quantum chemistry code.35 Convergence tests for electronic friction calculations can be found in Section 1.6 in the Supporting Information. We explicitly calculate electronic friction for C and O atoms. For the Cu atoms, we use isotropic electronic friction and a diagonal friction tensor. This friction value is calculated for a single atom in bulk copper (see Section S1.7 in the Supporting Information). Electron-phonon coupling calculations are done using the PBE54 exchange-correlation functional, as has been done in previous simulations of CO on Cu substrates.21, 35 Details can be found in Section S1.7 of the Supporting Information. We note that employing the HSE06 functional for electronic friction calculations in systems of this size is currently infeasible given available high-performance computing resources. Vibrational relaxation rates are determined from the friction tensor and the wass-weighted vibrational eigenmodes, , at the relevant vibrational frequency :
4.5 Molecular dynamics
We perform molecular dynamics (MD) simulations using the atomic simulation environment (ASE).55 A time step of is employed. Since we are interested in the dynamics of a single CO molecule, we conduct MD simulations in an extended unit cell containing a single adsorbed CO molecule, shown in Figure 1. The extended unit cell facilitates the back-folding of substrate phonon bands, enabling effective sampling of the metal phonon band structure.
To account for electron-phonon coupling, we use molecular dynamics with electronic friction, where the dynamics are governed by the Langevin equation:56
Here, represents the potential energy surface, which depends on the atomic positions . The term denotes the electronic friction tensor, while is the random force responsible for maintaining detailed balance at a given electronic temperature , as dictated by the second fluctuation-dissipation theorem.57 For the dynamics, the electronic friction tensor is calculated in the quasi-static limit with a perturbation frequency of zero. A comparison of lifetimes with and without a perturbing frequency can be found in Section S3.1 of the Supporting Information. During the calculations, the constant friction approximation is used for the electronic friction tensor, where the friction tensor values have been determined at the relaxed equilibrium geometry of the molecule on the surface.
To determine relaxation rates and lifetimes from the MD and MDEF trajectories, we use three different methods:
Equilibrium correlation analysis.
For each system, we first perform equilibration at , followed by computing a single, long MD or MDEF trajectory (see Section S3.2 of the Supporting Information). To analyse the velocity time series, we use normal mode decomposition and determine the cross-spectrum using the Welch method. Thereby, the velocity time series is cut into multiple, partially overlapping segments. A window function—a Hann window in our case—is then applied to each segment before the cross-spectrum is calculated. The cross-spectra of the different segments are then averaged. This improves the smoothness of the cross-spectrum. The height and width of the peak(s) in the cross-spectrum are then determined by fitting a Lorentzian function (see Section 4 of the Supporting Information).
Kinetic energy decay.
For each system and vibrational mode, relaxation rates and lifetimes are extracted from ensemble averages over 64 MD or MDEF trajectories. We first equilibrate the system at . Then, we excite the vibrational mode of interest by a displacement amplitude equivalent to along the respective eigenvector. Using normal mode decomposition, we project the velocity time series onto the basis of the vibration modes. We subsequently monitor the temporal decay of the envelope of the ensemble-averaged kinetic energy. Assuming exponential decay—as would be the case for a damped harmonic oscillator—we determine relaxation rates via an exponential fit.
Additive relaxation rates.
Total relaxation rates are calculated by summing the individual relaxation rates associated with phonon-phonon and electron-phonon coupling. Phonon-phonon relaxation rates are derived from the inverse values listed in Table 2, which were determined through equilibrium correlation analysis. Electron-phonon relaxation rates are obtained from the inverse values in Table 3, which were directly calculated from the electronic friction tensor.
5 Acknowledgments
The authors acknowledge financial support from the UKRI Future Leaders Fellowship program (MR/X023109/1), a UKRI Horizon grant (ERC StG, EP/X014088/1), and a UKRI Horizon grant (MSCA, EP/Y024923/1). We gratefully acknowledge valuable discussions with Nils Hertl and Connor Box. High-performance computing resources were provided via the Scientific Computing Research Technology Platform of the University of Warwick, the EPSRC-funded Materials Chemistry Consortium (EP/R029431/1, EP/X035859/1) and the UK Car-Parrinello consortium (EP/X035891/1) for the ARCHER2 UK National Supercomputing Service, the EPSRC-funded HPC Midlands+ computing centre for access to Sulis (EP/P020232/1), and EuroHPC Joint Undertaking and LUMI Consortium for access to LUMI (EHPC-EXT-2024E02-098).
6 Data availability
The data supporting this research has been uploaded to the NOMAD database at 10.17172/NOMAD/2025.04.29-1.